# OC-SOP: Enhancing Vision-Based 3D Semantic Occupancy Prediction by Object-Centric Awareness

### v1

@@ Abstract
Autonomous driving perception faces significant challenges due to occlusions and incomplete scene data in the environment. To overcome these issues, the task of semantic occupancy prediction (SOP) is proposed, which aims to jointly infer both the geometry and semantic labels of a scene from images. However, conventional camera-based methods typically treat all categories equally and primarily rely on local features, leading to suboptimal predictions—especially for dynamic foreground objects. To address this, we propose Object-Centric SOP (OC-SOP), a framework that integrates high-level object-centric cues extracted via a detection branch into the semantic occupancy prediction pipeline. This object-centric integration significantly enhances the prediction accuracy for foreground objects and achieves state-of-the-art performance among all categories on SemanticKITTI.
@@ II-A Camera-based 3D Reconstruction
Camera-based 3D perception aligns with the natural human ability to form a holistic scene understanding from rich visual information. It is also increasingly favored in autonomous driving due to its cost-effectiveness. Early multi-view reconstruction [5, 6] and SLAM [7, 8] studies estimated 3D geometry from corresponding 2D feature points using explicit mathematical constraints. However, their reliance on multi-view observations restricted agent movement and limited scene reconstruction quality. Eigen et al. [9] first introduced end-to-end monocular depth estimation using a CNN, enabling 3D reconstruction from a single 2D image. This addressed the ill-posed problem of 2D-to-3D reconstruction. Inspired by [9], some works focus on object-level single-view 3D reconstruction [10, 11, 12, 13], typically employing an encoder-decoder structure to learn explicit or implicit geometry representations. Dahnert et al. [14] expands this to scene-level reconstruction by lifting 2D panoptic features into 3D space. However, since single-view reconstruction focuses only on visible regions, it exhibits weak performance in predicting occluded areas, significantly hindering holistic 3D scene understanding. To address this, scene completion is proposed to predict unseen geometry based on observable information.
@@ II-B Semantic Occupancy Prediction (SOP)
Although there is some debate within the community regarding the definition of Semantic Occupancy Prediction (SOP), the widely accepted view is that SOP aims to predict occupancy status and semantics within a grid-based scene volume, covering both visible and occluded areas [15]. Based on the input modality, SOP can be categorized into LiDAR-based SOP [16, 17], Camera-based SOP [2], and Fusion-based SOP [18], among others. A broader task is Semantic Scene Completion (SSC), which jointly predicts geometry and semantics for both visible and occluded regions [19, 20]. However, scene representation is not limited to grid-based formats; it also includes explicit representations like point clouds [21] and meshes [22], as well as implicit representations such as Neural Radiance Fields (NeRF) [23]. Voxel-based representations are an effective way to model 3D scenes, as they discretize space into voxels, each storing attributes like occupancy, semantics, or learned feature vectors. Due to their compatibility with 3D convolutional operations, early LiDAR-based SOP [16, 17] methods commonly used U-Net architectures to predict occluded regions. Later, the 3D completion U-Net was extended to camera-based approaches. MonoScene [2] backprojects image features along optical rays to construct an initial voxel representation, which is then refined using a 3D completion U-Net. VoxFormer [24] introduced a Transformer-based approach, encoding spatial information as queries and context as key-value pairs, leveraging deformable attention for 3D-SOP.
@@ II-C Object-Centric Perception
Human perception is fundamentally object-centric, as our cognitive system naturally focuses on distinct objects, which forms the basis for advanced cognition and effective interaction with the world. Recent works have exploited this principle in both 3D object detection and video prediction. For instance, inspired by the end-to-end paradigm of DETR [25], DETR3D [26] establishes learnable 3D object queries that link 2D images via camera projection matrices, while OCVP [27] decomposes video frames into object components and models their dynamics and interactions to generate future video frames. Leveraging object-centric awareness for semantic occupancy prediction is a novel direction. Unlike object detection—which outputs coarse bounding box representations—semantic occupancy prediction (SOP) produces finer, voxel-level semantic maps. Moreover, compared to video prediction, which typically focuses on visible regions, SOP must also predict occupancy in unseen or occluded areas, making the incorporation of object-centric awareness even more critical for accurate scene understanding.
@@ III Method
As shown in Fig. 2, our network architecture comprises a main prediction branch and an object detection-based branch, which share a common feature lifting backbone. In the backbone, we adopt an encoder with dual decoders (EDD), where one decoder estimates depth and the other computes semantic content features. The semantic content features, combined with the estimated depth, are then fed into a feature lifting module that maps the semantic features into 3D space.
@@ IV-A1 Datasets
We train our model using two datasets. For voxel-wise prediction, we employ the SemanticKITTI [30] dataset, on which numerous classic baselines have been validated. Built on the KITTI odometry benchmark, SemanticKITTI enables to study semantic segmentation through manual annotations on raw LiDAR point clouds. The semantic scene completion benchmark is constructed by accumulating annotated scans within sequences and voxelizing them to form a semantic scene volume with dimensions of 256$\times$256$\times$32 (with each voxel representing 0.2 m). The dataset contains 21 classes, including 19 semantic labels, one free label, and one unknown label. In our experiments, we use the RGB images from cam-2 with a resolution of 1226$\times$370 (cropped to 1220$\times$370) and follow the official training/validation split of 3834 and 815 samples, respectively.
@@ V Conclusions and Outlook
We propose OC-SOP, which innovatively integrates object-centric awareness into the semantic occupancy prediction task. Our results on the SemanticKITTI test set demonstrate that OC-SOP significantly improves the accuracy of foreground object predictions. By focusing on other traffic participants, our approach enhances the perception system’s ability to understand complex traffic scenarios, ultimately contributing to safer decision-making in autonomous driving. In future work, we plan to explore additional ways of incorporating object-centric awareness into scene understanding, such as through Slot Attention [33].

### v2

@@ Abstract
Autonomous driving perception faces significant challenges due to occlusions and incomplete scene data in the environment. To overcome these issues, the task of semantic occupancy prediction (SOP) is proposed, which aims to jointly infer both the geometry and semantic labels of a scene from images. However, conventional camera-based methods typically treat all categories equally and primarily rely on local features, leading to suboptimal predictions, especially for dynamic foreground objects. To address this, we propose Object-Centric SOP (OC-SOP), a framework that integrates high-level object-centric cues extracted via a detection branch into the semantic occupancy prediction pipeline. This object-centric integration significantly enhances the prediction accuracy for foreground objects and achieves state-of-the-art performance among all categories on SemanticKITTI.
@@ II-A Camera-based 3D Reconstruction
Camera-based 3D perception aligns with the natural human ability to form a holistic scene understanding from rich visual information. It is also increasingly favored in autonomous driving due to its cost-effectiveness. Early multi-view reconstruction [5, 6, 7, 8] and SLAM [9, 10] studies estimated 3D geometry from corresponding 2D feature points using explicit mathematical constraints. However, their reliance on multi-view observations restricted agent movement and limited scene reconstruction quality. Eigen et al. [11] first introduced end-to-end monocular depth estimation using a CNN, enabling 3D reconstruction from a single 2D image. This addressed the ill-posed problem of 2D-to-3D reconstruction. Inspired by [11], some works focus on object-level single-view 3D reconstruction [12, 13, 14, 15], typically employing an encoder-decoder structure to learn explicit or implicit geometry representations. Dahnert et al. [16] expands this to scene-level reconstruction by lifting 2D panoptic features into 3D space. However, since single-view reconstruction focuses only on visible regions, it exhibits weak performance in predicting occluded areas, significantly hindering holistic 3D scene understanding. To address this, scene completion is proposed to predict unseen geometry based on observable information.
@@ II-B Semantic Occupancy Prediction (SOP)
Although there is some debate within the community regarding the definition of Semantic Occupancy Prediction (SOP), the widely accepted view is that SOP aims to predict occupancy status and semantics within a grid-based scene volume, covering both visible and occluded areas [17]. Based on the input modality, SOP can be categorized into LiDAR-based SOP [18, 19, 20], Camera-based SOP [2], and Fusion-based SOP [21], among others. A broader task is Semantic Scene Completion (SSC), which jointly predicts geometry and semantics for both visible and occluded regions [22, 23]. However, scene representation is not limited to grid-based formats; it also includes explicit representations like point clouds [24] and meshes [25], as well as implicit representations such as Neural Radiance Fields (NeRF) [26]. Voxel-based representations are an effective way to model 3D scenes, as they discretize space into voxels, each storing attributes like occupancy, semantics, or learned feature vectors. Due to their compatibility with 3D convolutional operations, early LiDAR-based SOP [18, 19] methods commonly used U-Net architectures to predict occluded regions. Later, the 3D completion U-Net was extended to camera-based approaches. MonoScene [2] backprojects image features along optical rays to construct an initial voxel representation, which is then refined using a 3D completion U-Net. VoxFormer [27] introduced a Transformer-based approach, encoding spatial information as queries and context as key-value pairs, leveraging deformable attention for 3D-SOP.
@@ II-C Object-Centric Perception
Human perception is fundamentally object-centric, as our cognitive system naturally focuses on distinct objects, which forms the basis for advanced cognition and effective interaction with the world. Recent works have exploited this principle in both 3D object detection [28] and video prediction. For instance, inspired by the end-to-end paradigm of DETR [29], DETR3D [30] establishes learnable 3D object queries that link 2D images via camera projection matrices, while OCVP [31] decomposes video frames into object components and models their dynamics and interactions to generate future video frames. Leveraging object-centric awareness for semantic occupancy prediction is a novel direction. Unlike object detection—which outputs coarse bounding box representations—semantic occupancy prediction (SOP) produces finer, voxel-level semantic maps. Moreover, compared to video prediction, which typically focuses on visible regions, SOP must also predict occupancy in unseen or occluded areas, making the incorporation of object-centric awareness even more critical for accurate scene understanding.
@@ III Method
As shown in Fig. 2, our dual-branch network architecture [32] comprises a main prediction branch and an object detection-based branch, which share a common feature lifting backbone. In the backbone, we adopt an encoder with dual decoders (EDD), where one decoder estimates depth and the other computes semantic content features. The semantic content features, combined with the estimated depth, are then fed into a feature lifting module that maps the semantic features into 3D space.
@@ IV-A1 Datasets
[…]
@@ V Conclusions and Outlook
We propose OC-SOP, which innovatively integrates object-centric awareness into the semantic occupancy prediction task. Our results on the SemanticKITTI test set demonstrate that OC-SOP significantly improves the accuracy of foreground object predictions. By focusing on other traffic participants, our approach enhances the perception system’s ability to understand complex traffic scenarios, ultimately contributing to safer decision-making in autonomous driving. In future work, we plan to explore emerging techniques in prompt-driven representation learning [38, 39, 40] and multimodal fusion [41, 42] to further improve object-centric semantics and generalization across diverse sensing conditions.